Write 116 2/3% as a fraction or a mixed number. OK I asked this question before and got the same answer from 2 different people AND I also thought the same thing. BUT 350/3 is not an option. This is why i am asking the ? Here are the answers that are available:
A) 1/6
B) 1 1/6
C) 16 2/3
D) 1 5/6
Help me please.


Answer by stanbon(75887)

Write 116 2/3% as a fraction or a mixed number. OK I asked this question before and got the same answer from 2 different people AND I also thought the same thing. BUT 350/3 is not an option. This is why i am asking the ? Here are the answers that are available:
A) 1/6
B) 1 1/6
C) 16 2/3
D) 1 5/6
----------------
116 2/3% = 116.6666/100 = 1.1666666..
= 1.1667
= 1 + 0.1667
= 1 1/6
=================
Cheers,
Stan H.